OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 5 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need?

hartnn (hartnn):

if this question is to find area of curved portion of cylinder,then the formula is 2*pi*r*h =2*pi*5*6 =60pi =188.49 square inches